OpenStudy (anonymous):

Def Eq Solve: (2xy+x)dx+(x^2+y)dy=0

OpenStudy (anonymous):

2xy dx+x^2 dy + xdx+ydy=0 Observe that dervative of x^2 y=2xydx+x^2dy(The first two terms) d(x^2 y) + xdx+ydy=0 integrate x^2y+(x^2+y^2)/2=0

OpenStudy (anonymous):

Omg why are you so fast at this xD Im still pretty lost though on this

OpenStudy (anonymous):

lol. integral of dx=x so integral of d(x^2 y) is also x^2 y.

OpenStudy (anonymous):

well not exactly this problem but other ones where you may not be able to observe that? or will there always be something like that

OpenStudy (anonymous):

Well there are some standard stuff like xdy-ydx/x^2=d(x/y) and all.

OpenStudy (anonymous):

yea is there another way you could possibly show it assuming I will most likely not see that while I'm doing my test xD

OpenStudy (anonymous):

Just learn the standard results for the test its not very hard to identify, all equation are always solvable.

OpenStudy (anonymous):

I can do that a lot with calc but for some reason DE is just not sinking in at all... lol